OpenStudy (anonymous):

What is the measure of the smaller angle formed by the hour and minute hands at 1:28? please show me the steps.

OpenStudy (campbell_st):

on a clock each minute = 6 degrees.... but is the hour hand still on 1o'clock or has it moved to some point between 1 and 2...? In real life the hour hand would have moved...

OpenStudy (anonymous):

@SAMMY, you actually posted this kind of problem, yesterday. . am i right?

OpenStudy (anonymous):

yes.but it's little different from that

OpenStudy (dumbcow):

yes the hour hand would have moved proportionately so it would be 28/60 hrs past 1 there are 30 degrees in between each hour so hour hand is 30+30(28/60) = 44 degrees past 12 minute hand is 6*28 = 168 degrees past 12 --> 168 - 44 = 124 degrees

OpenStudy (anonymous):

Thanks a lot for solving it

OpenStudy (anonymous):

{28- [(28/60)(5)+5]} (6)= 124 degrees.